# Higher Memory Effects in Numerical Simulations of Binary Black Hole Mergers

### v2

@@ I Introduction
The “displacement” gravitational wave memory effect arises as a permanent change in the separation of two observers due to the passage of gravitational waves. This effect was first postulated in the linearized regime Zel’dovich and Polnarev (1974), but for bound systems, the predominant effect is sourced by nonlinear effects in the propagation of gravitational waves Christodoulou (1991); Thorne (1992). This effect is difficult to detect both because of its low-frequency nature and that it is simply smaller in amplitude than the predominantly oscillatory part of the gravitational wave signal. However, there have been numerous proposals to measure it using both ground-based detectors (by stacking multiple signals in current detectors Lasky et al. (2016); Boersma et al. (2020); Grant and Nichols (2023), which has been performed explicitly in Refs. Hübner et al. (2020, 2021)) or from single events in future detectors Johnson et al. (2019); Grant and Nichols (2023). It should also be seen by space-based detectors Islo et al. (2019); Gasparotto et al. (2023); Ghosh et al. (2023) and pulsar timing arrays, and the latter currently provide upper bounds on the memory coming from the stochastic background Agazie et al. (2023).
@@
In addition to the displacement memory, a variety of other effects appearing as permanent changes in some idealized detector have been considered. These effects all fall under the class of so-called “persistent observables” Flanagan et al. (2019), the simplest subset of which forms the so-called “curve deviation”. In Ref. Grant and Nichols (2022), it was found that the curve deviation could be written in terms of a set of quantities at null infinity that are temporal moments [^1] These temporal moments can be related to the integer values of the Mellin transform: see the discussion in Sec. II.4 below. of the Bondi news: a tensor which characterizes the presence of radiation at null infinity. For the zeroth moment (that is, a single integral) of the news, this relationship was already known, and reflects the fact that integrating the news between early and late times gives the value of the displacement memory effect. However, the curve deviation observable depends on higher moments of the news as well. For example, the “drift” memory (previously called “subleading displacement memory” Flanagan et al. (2019): see Ref. Grant (2023) for an explanation for the change in terminology), which described the separation of two observers with a nonzero initial relative velocity, was related to the first and zeroth moments of the news. Similarly, higher moments of the news could be used to determine the dependence of the final separation on the initial relative acceleration and higher derivatives. Collectively, we refer to these higher moments of the news as “higher memory effects”.
@@
Furthermore, an important result of Ref. Grant and Nichols (2022) was that these moments of the news could be written in terms of a “charge” contribution and a “flux” contribution, which generalizes the usual linear/nonlinear or ordinary/null Bieri and Garfinkle (2014) splittings of the displacement memory. For the first moment, such a decomposition was already known as well, as it was shown for spin Pasterski et al. (2016); Nichols (2017) and center-of-mass Nichols (2018) memories, which form the magnetic and electric parts of the drift memory. That such a splitting could be achieved for all higher memory effects, however, was a novel result, and follows from the asymptotic form of the Einstein equations.
@@
There is another, useful formulation of the displacement memory in terms of asymptotic symmetries. Here, for each spherical harmonic mode, the charge contribution can be thought of as the change in a charge constructed from a member of the supertranslation subalgebra of the Bondi-Metzner-Sachs (BMS) algebra—the symmetry algebra at null infinity. Understanding the drift memory in this way, however, requires an extension of the BMS algebra into either of the “extended” Barnich and Troessaert (2010) or “generalized” Campiglia and Laddha (2014) BMS algebras, and many issues with these extensions have been pointed out in the literature Flanagan and Nichols (2017); Flanagan et al. (2020); Elhashash and Nichols (2021). Recently, an infinite set of charges were also defined which form a representation of the algebra $Lw_{1+\infty}$ Freidel et al. (2022a), and at linear order these charges can be shown Compère et al. (2022) to be equivalent to those in Ref. Grant and Nichols (2022). While these connections to symmetry algebras are interesting, we do not discuss them further in this paper, as everything that we present can be straightforwardly derived from the Einstein equations.
@@
An important question remains about the status of this split into charge and flux contributions: for a given system, which is larger? This is important for observing the memory effect, by the following argument Grant and Nichols (2023). For simplicity, we restrict to a discussion of the displacement memory, although similar arguments for higher moments also apply. First, note that by “detecting the displacement memory”, one does not mean “measuring some finite offset in the detector”, as current ground-based detectors are only sensitive in some frequency range that does not include zero. Instead, one should try to detect a part of the signal which contributes to this finite offset, which one could call the time-dependent “memory signal”. In the case of the displacement memory, it happens to be the case that, for binary mergers, it is the flux contribution which makes the largest contribution (for the $m=0$ modes) to the zeroth moment of the news (see, for example, Ref. Mitman et al. (2020)). Moreover, since this contribution as a function of time is non-oscillatory and looks like a smoothed out step function, it is a reasonable choice for the time-dependent “displacement memory signal”.
@@
When considering higher moments of the news, one is faced with another issue: current ground-based gravitational wave detectors are sensitive to the shear, and not any particular integrals of the news. If one is given the $n$th moment of the news as a function of time, one can obtain the shear as a function of time (up to an arbitrary constant) by differentiating $n$ times. As such, the question of which is the larger contribution should be asked of the $n$th derivative of the charge or flux. For the case of the spin memory (the magnetic part of the first moment of the news), it is once again the flux contribution that is larger. Furthermore, this contribution is again non-oscillatory Mitman et al. (2020) (in the $m=0$ modes) and resembles a single bump, i.e., a smoothed-out delta function.
@@
    Which contribution to the shear, charge or flux, is more important?
@@
As there are an infinite number of these higher memory effects, we restrict our attention to the displacement memory (zeroth moment), drift memory (first moment), and the second moment of the news. To explore these inquiries, we will examine the waveforms produced by numerical relativity simulations of binary black hole mergers that have been run using the Simulating eXtreme Spacetimes (SXS) Collaboration’s SpEC code (to perform the Cauchy evolution) and SpECTRE code [to extract the asymptotic data via a Cauchy-characteristic evolution (CCE)] Deppe et al. (2023); Moxon et al. (2020, 2023). The code which we used to produce these waveforms is publicly available at Mitman (2024).
@@
In this paper, we set $G=c=1$ and adopt the mostly plus metric signature. We use Latin characters from the beginning of the alphabet ($a$, $b$, etc.) for abstract indices and Latin characters from the middle of the alphabet ($i$, $j$, etc.) in order to denote angular indices on the two-sphere.
@@ II.2 GHP formalism
Note that, if $v$ has spin weight $s$, then ð$\eth v$ has spin weight $s+1$ and ð¯$\bar{\eth}v$ has spin weight $s-1$. As such, we can consider the eigenfunctions of $\eth\bar{\eth}$ (or, equivalently, $\bar{\eth}\eth$), which are the spin-weighted spherical harmonics ${}_{s}Y_{\ell m}$. In this paper, we will only need the following properties (using the conventions of the scri package Boyle et al. (2023)):
@@ II.4 Definition of our observables
Here, we explicitly keep the “reference time” $\tilde{u}$ unspecified. In Ref. Grant and Nichols (2022), these moments of the news could be related to pieces of the curve deviation observable if $\tilde{u}=u_{0}$. We call such moments the “Mellin moments”, since (for $u_{0}=0$ and $u_{1}=\infty$), these moments are related to the Mellin transform $\mathcal{M}$ of $\dot{\bar{\sigma}}$, which is defined by
$\mathcal{M}_{n}\{f\}\equiv\int_{0}^{\infty}u^{n-1}f(u)\mathrm{d}u$ (28)
(for some arbitrary function $f$). The exact relation is
@@
where the factor of $(-1)^{n}$ was absorbed into the definition of $\mathcal{N}_{n}$ in Ref. Grant and Nichols (2022). This interpretation in terms of the Mellin transform arises in discussions of celestial holography Freidel et al. (2022b).
@@
In this section, we review the results of Ref. Grant and Nichols (2022), which showed that one could write these moments of the news $\mathcal{N}_{n}(u_{1},u_{0};\tilde{u})$ in terms of two contributions: a change in a “charge” and an integral of a “flux”. This is analogous to the split into “linear” and “nonlinear” memory of Ref. Zel’dovich and Polnarev (1974) and Ref. Christodoulou (1991), or of “ordinary” and “null” of Ref. Bieri and Garfinkle (2014), in the case where one is considering vacuum general relativity. [^3] In the case where there exist null matter fields, for example in Einstein-Maxwell theory, the stress-energy tensor of these fields is included in the “null” part defined by Ref. Bieri and Garfinkle (2014), whereas here we would place it in a third category in this splitting.
The general definition of a “charge” which we use here is that a charge is a quantity, such as the mass aspect $m$, which is constant in regions where the news vanishes. In this sense, these are quasi-conserved quantities near null infinity. There are a variety of prescriptions to define charges near null infinity, in the sense of quasi-conserved quantities which are conjugate to some symmetry Wald and Zoupas (2000); Compère et al. (2020); Grant et al. (2022); Elhashash and Nichols (2021); however, following Ref. Grant and Nichols (2022), we merely provide a procedure by which one could construct these charges, and do not consider their relationship to (potential) symmetries at null infinity.
@@ III.1 First three moments of the news
Note that Eq. (32) does not give the entire zeroth moment of the news: we are missing $\operatorname{Im}[\eth^{2}\mathcal{N}_{0}(u_{1},u_{0};\tilde{u})]$, and there is no equation analogous to Eq. (32) from which this quantity can be derived. For reasons discussed further in Refs. Flanagan and Nichols (2017); Mädler and Winicour (2016), it is expected to be small, which is consistent with Fig. 4, as well as the results of Sec. IIID of Ref. Mitman et al. (2020).
@@
and
$\mathcal{F}_{1}\equiv-\dot{\bar{\sigma}}\eth\sigma-3\sigma\eth\dot{\bar{\sigma}}.$ (38)
Integrating as before, we find that
@@
Again, this quantity is a charge: since
@@ III.2 Ambiguities
Note that there exist non-trivial modifications of the “charges” $m$, $\tilde{\psi}_{1}$, and $\tilde{\psi}_{0}$, that still satisfy the requirements given above. First, by “non-trivial”, we mean that these modifications modify the charges in regions where $\dot{\sigma}=0$. As such, they can only depend on $\tilde{\psi}_{2}$, $\tilde{\psi}_{1}$, $\tilde{\psi}_{0}$, and $\sigma$; these are the only choices which satisfy the requirement that their derivatives with respect to time vanish when $\dot{\sigma}=0$.
@@
To allow for easier comparison to the literature, we adopt the Wald-Zoupas convention for the charges and fluxes for the remainder of the paper.
@@ IV.1 Numerical Data
We stress that the numerical error estimate that we provide is more an error estimate on the lower-resolution simulation, and that the errors for the simulation that we examine will be smaller than what is shown. With this in mind, we find that the violation of the Bianchi identities and the numerical error for the Weyl scalars $\psi_{2}$, $\psi_{3}$, and $\psi_{4}$ are particularly low with absolute errors of $\lesssim 0.001\%$, whereas for the Weyl scalars $\psi_{0}$ and $\psi_{1}$ the Bianchi violations are noticeably larger. We attribute this to the fact that both $\psi_{0}$ and $\psi_{1}$ are more influenced by backscattering physics, that is, the fact that information should be traveling back and forth between the Cauchy evolution and CCE, than the other Weyl scalars. In fact, this notion has even been confirmed by the recent implementation of SpECTRE’s Cauchy-characteristic matching code, which shows the violation of the $\psi_{0}$ and $\psi_{1}$ Bianchi identities to be much smaller than what CCE usually yields Ma et al. (2023). [^4] In particular, see Fig. 7 of Ref. Ma et al. (2023). Nonetheless, because the violations of the $\psi_{0}$ and $\psi_{1}$ Bianchi identities still correspond to absolute errors of $\lesssim 1\%$, the results of Fig. 1 indicate that the errors on our asymptotic data are reasonable enough to perform the following analysis. Furthermore, we find that the results drawn from the following analysis are unchanged if one uses the lower-resolution simulation instead of the higher-resolution simulation.

### v3

@@ I Introduction
The gravitational wave memory effect (now called “displacement memory” to avoid confusion with other, similar effects) arises as a permanent change in the separation of two observers due to the passage of gravitational waves. This effect was first postulated in the linearized regime Zel’dovich and Polnarev (1974), but for bound systems, the predominant effect is sourced by nonlinear effects in the propagation of gravitational waves Christodoulou (1991); Thorne (1992). This effect is difficult to detect both because of its low-frequency nature and that it is simply smaller in amplitude than the predominantly oscillatory part of the gravitational wave signal. However, there have been numerous proposals to measure it using both ground-based detectors (by stacking multiple signals in current detectors Lasky et al. (2016); Boersma et al. (2020); Grant and Nichols (2023), which has been performed explicitly in Refs. Hübner et al. (2020, 2021)) or from single events in future detectors Johnson et al. (2019); Grant and Nichols (2023). It should also be seen by space-based detectors Islo et al. (2019); Gasparotto et al. (2023); Ghosh et al. (2023) and pulsar timing arrays, and the latter currently provide upper bounds on the memory coming from the stochastic background Agazie et al. (2024).
@@
In addition to the displacement memory, a variety of other effects have been considered in the literature which appear as permanent changes in some idealized detector (for a summary, see Flanagan et al. (2019)). While this paper does not focus on the observable nature of these effects, the “curve deviation” observable of Flanagan et al. (2019) provides a motivation for studying the quantities of interest in this paper as a simple generalization of the displacement memory effect. As shown in Grant and Nichols (2022), the curve deviation near null infinity can be written in terms of the temporal moments [^1] These temporal moments can be related to the integer values of the Mellin transform: see the discussion in Sec. II.4 below. of the Bondi news, a tensor which characterizes the presence of radiation at null infinity. Much like the moments of a probability distribution, these moments characterize the Bondi news as a function of time. The zeroth moment (that is, a single integral) of the news is exactly the displacement memory effect, but with each higher moment of the news one can recover additional observables. For example, the spin Pasterski et al. (2016); Nichols (2017) and center-of-mass Nichols (2018) memories form the electric- and magnetic-parity pieces of the “drift memory” [^2] Note that this effect was previous called the “subleading displacement memory” in Flanagan et al. (2019); see Ref. Grant (2024) for an explanation for the change in terminology., which is related to the first moment of the news. Collectively, we refer to the infinite collection of moments of the news as “higher memory effects”.
@@
More relevantly for this paper, Ref. Grant and Nichols (2022) showed that these moments of the news could be written in terms of a “charge” contribution and a “flux” contribution, which generalizes the usual linear/nonlinear or ordinary/null Bieri and Garfinkle (2014) splittings of the displacement memory. For the drift memory, such a decomposition was already known Pasterski et al. (2016); Nichols (2018), but that such a splitting could be achieved for all higher memory effects was a novel result, and follows from the asymptotic form of the Einstein equations. We review this construction explicitly for the zeroth, first, and second moments in Sec. III.1 below.
@@
Asymptotic symmetries provide another formulation of the displacement memory that is frequently discussed in the literature (see, for example, Strominger (2018) and the references therein). Here, for each spherical harmonic mode, the charge contribution can be thought of as the change in a charge constructed from a member of the supertranslation subalgebra of the Bondi-Metzner-Sachs (BMS) algebra—the symmetry algebra at null infinity. Understanding the drift memory in this way, however, requires an extension of the BMS algebra into either of the “extended” Barnich and Troessaert (2010) or “generalized” Campiglia and Laddha (2014) BMS algebras, and many issues with these extensions have been pointed out in the literature Flanagan and Nichols (2017); Flanagan et al. (2020); Elhashash and Nichols (2021). Recently, an infinite set of charges were also defined which form a representation of the algebra $Lw_{1+\infty}$ Freidel et al. (2022), and at linear order these charges can be shown Compère et al. (2022) to be equivalent to those in Ref. Grant and Nichols (2022). While these connections to symmetry algebras are interesting, we do not discuss them further in this paper, as everything that we present can be straightforwardly derived from the Einstein equations.
@@
An important question remains about the status of this split into charge and flux contributions: for a given system, which is larger? This is important for observing the memory effect, by the following argument Grant and Nichols (2023). For simplicity, we restrict to a discussion of the displacement memory, although similar arguments for higher memories also apply. First, note that by “detecting the displacement memory”, one does not mean “measuring some finite offset in the detector”, as current ground-based detectors are only sensitive in some frequency range that does not include zero. Instead, one should try to detect a part of the signal which contributes to this finite offset, which one could call the time-dependent “memory signal”. In the case of the displacement memory, it happens to be the case that, for binary mergers, it is the flux contribution which makes the largest contribution (for the $m=0$ modes) to the zeroth moment of the news (see, for example, Ref. Mitman et al. (2020)). Moreover, since this contribution as a function of time is non-oscillatory and looks like a smoothed out step function, it is a reasonable choice for the time-dependent “displacement memory signal”.
@@
When considering higher memories, one is faced with another issue: current ground-based gravitational wave detectors are sensitive to the shear, and not any particular integrals of the news. If one is given the $n$th moment of the news as a function of time, one can obtain the shear as a function of time (up to an arbitrary initial value) by differentiating $n$ times. [^3] See Eq. (27) for the explicit relationship between the shear and the moments of the news. As such, the question of which is the larger contribution should be asked of the $n$th derivative of the charge or flux. For the case of the spin memory (the magnetic part of the first moment of the news), it is once again the flux contribution that is larger. Furthermore, this contribution is again non-oscillatory Mitman et al. (2020) (in the $m=0$ modes) and resembles a single bump, i.e., a smoothed-out delta function.
@@
    Which quantity contributes more significantly to the shear: charge or flux?
@@
As there are an infinite number of these higher memory effects, we restrict our attention to those associated with the zeroth, first, and second moments of the news: the displacement, drift, and “ballistic” Grant (2024) memory effects. To explore these questions, we will examine the waveforms produced by numerical relativity simulations of binary black hole mergers that have been run using the Simulating eXtreme Spacetimes (SXS) Collaboration’s SpEC code (to perform the Cauchy evolution) and SpECTRE code [to extract the asymptotic data via a Cauchy-characteristic evolution (CCE)] Deppe et al. (2023); Moxon et al. (2020, 2023). The code which we used to produce these waveforms is publicly available at Mitman (2024).
@@
In this paper, we set $G=c=1$ and adopt the mostly plus metric signature. We use Latin characters from the beginning of the alphabet ($a$, $b$, etc.) for abstract indices and Latin characters from the middle of the alphabet ($i$, $j$, etc.) in order to denote indices associated with coordinates on the two-sphere (for example, spherical coordinates $\{\theta,\phi\}$).
@@ II.2 GHP formalism
Note that, if $v$ has spin weight $s$, then ð$\eth v$ has spin weight $s+1$ and ð¯$\bar{\eth}v$ has spin weight $s-1$. As such, we can consider the eigenfunctions of $\eth\bar{\eth}$ (or, equivalently, $\bar{\eth}\eth$), which are the spin-weighted spherical harmonics ${}_{s}Y_{\ell m}$. In this paper, we will only need the following properties (using the conventions of the scri package Boyle et al. (2023) [^5] The scri package is a module for constructing, transforming, and working with waveform data at future null infinity Boyle et al. (2023).):
@@ II.4 Definition of our observables
Here, we explicitly keep the “reference time” $\tilde{u}$ unspecified. In Ref. Grant and Nichols (2022), these moments of the news could be related to pieces of the curve deviation observable if $\tilde{u}=u_{0}$. We call such moments the “Mellin moments”, since, for $u_{0}=0$ and $u_{1}=\infty$, these moments are related to the Mellin transform $\mathcal{M}$ of $\dot{\bar{\sigma}}$, which is defined by
$\mathcal{M}_{n}\{f\}\equiv\int_{0}^{\infty}u^{n-1}f(u)\mathrm{d}u,$ (28)
for some arbitrary function $f$. The exact relation is
@@
where the factor of $(-1)^{n}$ was absorbed into the definition of $\mathcal{N}_{n}$ in Ref. Grant and Nichols (2022). This interpretation in terms of the Mellin transform arises in discussions of celestial holography Freidel et al. (2024).
@@
In this section, we review the results of Ref. Grant and Nichols (2022), which showed that one could write these moments of the news $\mathcal{N}_{n}(u_{1},u_{0};\tilde{u})$ in terms of two contributions: a change in a “charge” and an integral of a “flux”. This is analogous to the split into “linear” and “nonlinear” memory of Ref. Zel’dovich and Polnarev (1974) and Ref. Christodoulou (1991), or of “ordinary” and “null” of Ref. Bieri and Garfinkle (2014), in the case of vacuum general relativity. [^6] In the case where there exist null matter fields, for example in Einstein-Maxwell theory, the stress-energy tensor of these fields is included in the “null” part defined by Ref. Bieri and Garfinkle (2014), whereas here we would place it in a third category in this splitting.
The general definition of a “charge” which we use here is that a charge is a quantity, such as the mass aspect $m$, which is constant in regions where the news vanishes. In this sense, these are quasi-conserved quantities near null infinity. There are a variety of prescriptions to define charges near null infinity, in the sense of quasi-conserved quantities which are conjugate to some symmetry Wald and Zoupas (2000); Compère et al. (2020); Grant et al. (2022); Elhashash and Nichols (2021); however, following Ref. Grant and Nichols (2022), we merely provide a procedure by which one could construct these charges, and do not consider their relationship to (potential) symmetries at null infinity. Note, however, that our charges for the zeroth and first moments of the news are consistent with those in previous works, like Wald and Zoupas (2000); Compère et al. (2020).
@@ III.1 First three moments of the news
Here Eq. (32) does not give the entire zeroth moment of the news: we are missing $\operatorname{Im}[\eth^{2}\mathcal{N}_{0}(u_{1},u_{0};\tilde{u})]$, and there is no equation analogous to Eq. (32) from which this quantity can be derived. For reasons discussed further in Refs. Flanagan and Nichols (2017); Mädler and Winicour (2016), it is expected to be fairly subdominant, which is consistent with Fig. 2 and Fig. 4.777This is primarily due to the fact that magnetic memory is not sourced in binary black hole mergers, and is only known to appear for specific classes of spacetimes Bieri (2021); Satishchandran and Wald (2019). Last, note that Eq. (32) can also be written without the ð$\eth^{2}$ operator by simply inverting this operation via Eq. (16).
@@
for
$\mathcal{F}_{1}\equiv-\dot{\bar{\sigma}}\eth\sigma-3\sigma\eth\dot{\bar{\sigma}},$ (38)
which implies that $\dot{\tilde{\psi}}_{1}=0$ when $\dot{\sigma}=0$. Integrating as before, we find that
@@
Again, this quantity is a charge, since
@@ III.2 Ambiguities
Note that there exist non-trivial modifications of the “charges” $m$, $\tilde{\psi}_{1}$, and $\tilde{\psi}_{0}$, that still satisfy the requirement that they are charges, namely that their time derivatives vanish in the absence of news. By “non-trivial”, we mean that the modifications do not vanish in regions where the news vanishes, which eliminates any modifications using $\psi_{3}$ or $\psi_{4}$, as they vanish in such regions. In order for their time derivatives to vanish in the absence of news, any modifications must be constructed from quantities which themselves have that property, which eliminates all quantities other than $\tilde{\psi}_{2}$, $\tilde{\psi}_{1}$, $\tilde{\psi}_{0}$, and $\sigma$.
@@
To allow for easier comparison to the literature, we adopt the Wald-Zoupas convention for the charges and fluxes for the remainder of the paper. Note, however, that only the sum of the charge and flux contributions is invariant, and so any claims about the relative magnitude the contributions, or their morphology, will be affected by this choice. In particular, the particular shape of the flux contribution to the center-of-mass memory, as given in the left panel of Fig. 3, will be different in the case where one makes another choice, although the discrepancy appears to be small and the qualitative features the same.
@@ IV.1 Numerical Data
We stress that the numerical error estimate that we provide is more an error estimate on the lower-resolution simulation, and that the errors for the simulation that we examine will be smaller than what is shown. With this in mind, we find that the violation of the Bianchi identities and the numerical error for the Weyl scalars $\psi_{2}$, $\psi_{3}$, and $\psi_{4}$ are particularly low with absolute errors of $\lesssim 0.001\%$, whereas for the Weyl scalars $\psi_{0}$ and $\psi_{1}$ the Bianchi violations are noticeably larger. We attribute this to the fact that both $\psi_{0}$ and $\psi_{1}$ are more influenced by backscattering physics, that is, the fact that information should be traveling back and forth between the Cauchy evolution and CCE, than the other Weyl scalars. In fact, this notion has even been confirmed by the recent implementation of SpECTRE’s Cauchy-characteristic matching code, which shows that the violation of the $\psi_{0}$ and $\psi_{1}$ Bianchi identities is much smaller than what CCE usually yields, at least for problem of evolving a Teukolsky wave Ma et al. (2024). [^8] In particular, see Fig. 7 of Ref. Ma et al. (2024). Nonetheless, because the violations of the $\psi_{0}$ and $\psi_{1}$ Bianchi identities still correspond to absolute errors of $\lesssim 1\%$, the results of Fig. 1 indicate that the errors on our asymptotic data are reasonable enough to perform the following analysis. Furthermore, we find that the results drawn from the following analysis are unchanged if one uses the lower-resolution simulation instead of the higher-resolution simulation.